### v1

Abstract
Forest fire prediction involves estimating the likelihood of fire ignition or related risk levels in a specific area over a defined time period. With climate change intensifying fire behavior and frequency, accurate prediction has become one of the most pressing challenges in Artificial Intelligence (AI). Traditionally, fire ignition is approached as a binary classification task in the literature. However, this formulation oversimplifies the problem, especially from the perspective of end-users such as firefighters. In general, as is the case in France, firefighting units are organized by department, each with its terrain, climate conditions, and historical experience with fire events. Consequently, fire risk should be modeled in a way that is sensitive to local conditions and does not assume uniform risk across all regions. This paper proposes a new approach that tailors fire risk assessment to departmental contexts, offering more actionable and region-specific predictions for operational use. With this, we present the first national-scale AI benchmark for metropolitan France using state-of-the-art AI models on a relatively unexplored dataset. Finally, we offer a summary of important future works that should be taken into account. Supplementary materials are available on github.
1 Introduction
The immediate cost of wildfire suppression, compensation for damaged property, and the aftermath of lost agricultural productivity and tourism put a strain on local economies. For instance, the 2017 wildfires in the Mediterranean region resulted in estimated economic damages exceeding €200 million. From a human perspective, wildfires in France often lead to evacuations, sometimes involving thousands of residents and tourists. Health risks related to smoke inhalation and post-fire water contamination are also significant concerns [9, 3, 2]. Ecologically, these fires can cause irreversible harm to unique ecosystems. Additionally, wildfires contribute to topsoil erosion, threatening agriculture and leading to landslides in hilly terrains. Lastly, the emissions from these fires contribute not only to local air pollution but also to France’s national carbon footprint, further complicating efforts to combat climate change. According to reporterre, more than 4,000 hectares have burned in 2024.
2 State of the art
Constructing a wildfire prediction model has been central to wildfire research. We define the wildfire prediction in definition 1.
Definition 1
Wildfire prediction : Predict the future risk of wildfire ignition (or a linked-value) $V$ in a particular area $A$, for a particular time range $T$. Knowing a particular set of characteristics $F$, we can mathematically define the risk of wildfire as:
$V(A,T)=R(F(A,T))$
where R is a particular set of functions or an algorithm that calculates the risk value.
One of the most widely used wildfire risk assessment tools is the Fire Weather Index (FWI), developed in Canada in 1969. This system generates several fire risk indicators based on meteorological inputs such as temperature, precipitation, and wind speed. In France, the FWI has been adapted and studied for local contexts. Pimont et al. [15] employed Generalized Additive Models (GAMs) to estimate fire occurrence in Bouches-du-Rhône using FWI and forest cover data. Dupire et al. [8] analyzed the vulnerability of forest ecosystems to fire in the French Alps. Among these studies, the Mediterranean basin emerges as the most extensively examined region, due to its heightened susceptibility to wildfires.
Several studies worldwide have analyzed monthly forest fire probabilities using various methods. Zhang et al.[17] analyzed monthly forest fire probabilities in Changsha, China, using a Random Forest model. They identified vegetation as a key factor for these predictions. Their dataset (2004–2021) employed a 1:1 sampling ratio of fire to non-fire points, which introduces bias by excluding many potential non-fire locations. Chen et al.[5] also employed a GNN to predict fire size in Portugal, using the publicly available MontesinhoFires dataset. While this dataset includes features such as the Fire Weather Index (FWI) and burn duration, it covers only 2000–2003 and is therefore outdated. Monthly fire prediction has also been studied in Australia by Cisneros et al.[6]. The paper adopts a two-stage deep graphical regression: it first predicts the monthly probability that each Australian SA1/SA2 region experiences at least one wildfire, using a graph-convolutional neural network logistic model. Conditional on an ignition, it then fits an extended generalised Pareto distribution to the square‑root burned area, so the same framework quantifies both ignition risk and the expected severity of the area burned.
Mutakabbir et al. [14] train straightforward deep neural-network classifiers and regressors on data generated by Spatio-Temporal Agnostic Sampling (STAS), which balances each wildfire with a non-fire sequence taken M months earlier from the exact same sensor location so that the model learns feature changes rather than absolute spatio-temporal cues. All reported metrics come from an 80%/20% random train‑test split of the 1998‑2018 records—no dedicated temporal hold‑out was used in the main experiments, so training and test samples are intermixed in time
Weekly fire prediction has been attempted by Michail et al.[12]. They used GraphCast (Lam et al.[11]) coupled with a temporal encoder achieving a solid global AUPRC of 0.64 for one-period-ahead prediction. However, when the same model is evaluated per region, its score in Europe collapses to only $0.20$—one of the weakest of all regions, underlining that the overall metric masks large spatial disparities. This gap shows that a single ”global” network does not automatically generalize across continents and argues for tailoring or fine-tuning models to specific geographic domains, a need the authors themselves acknowledge. Additionally, the original GraphCast architecture was not tested.
Reducing wildfire danger to a binary outcome masks the strong regional heterogeneity of risk in France. Two ignitions in low-exposure Brittany can signal a crisis, while the same count in Mediterranean departments is routine. Standard models flatten these contrasts, trivializing near-certain July fires in the south and exaggerating sparse events elsewhere. Such uniform treatment misguides resource allocation because danger must be expressed relative to local climatology and capacity. Additionally, we found no analysis of any French department other than the Mediterranean basin, whereas, as shown in Figure 1, other regions have a non-negligible number of fires. This was particularly true in 2022 as discussed by Hardy Karim [10]. We therefore present the first national-scale AI benchmark for metropolitan France, replacing binary labels with a region-aware multi-class scheme. Our framework maps risk at the departmental level, directly aligning predictions with operational firefighting jurisdictions. Using a relatively unexplored dataset, we evaluate several state-of-the-art models across four forecast horizons and two targets. While the prediction of the highest classes is still an issue, the comparison between multi-class and binary variants shows gains in predicting fires in the highest risk regions, along with a better target interpretation. We also provide a data-driven strategy for selecting the best model for each horizon. These findings open new directions, from dynamic class adaptation to region-specific covariate design, for truly nuanced wildfire-risk prediction.
Figure 1: French forest fire distribution across all departments between 2017-06-12 and 2023-01-01. For visualisation purposes, we only selected departments where the total number of forest fires was greater than 50.
3 Constructing the database
Figure 2: Database construction process applied in this study. Apart from the target-related process, for which we show Bouches-du-Rhône, the department shown is Ain.
3.1 Processing fire occurrence and burned area
BDIFF is an internet application designed to centralize all data on forest fires across French territory since 2006 and make this information available to the public and state services. Fire locations are referenced by the name of the city where they occurred; coordinates are not provided. Fire are defined, notably, by the burned area.
We collected fire data from June 12, 2017, to December 31, 2023. We generated daily features for each French department during this period. The years 2017–2020 and 2022 were used for training, 2021 for validation, and 2023 for testing. Due to the severity of fire activity in 2022, it was included in the training set, while a more typical year (2021) was selected for validation.
City names were used to spatially locate fire points and generate 2 $\mathrm{km}\times 2\mathrm{~{}km}$ rasters per department. Two targets were defined: number of fires (FO) and burned (BA) area, across 4 time ranges: current day and the next 7, 15, and 31 days. For each day and department, we summed each target to obtain the total number of fires and burned area that occurred. We transformed the classical binary and regression problem into a multi-class classification problem. For each department, we created an ordinal 5-class signal for both occurrence and burned area using the K-Means algorithm. Each sample with 0 fires was set to class 0, and positive samples were clustered into 4 classes, representing None, Normal, Medium, High, and Extreme fire occurrence or burned area. With this representation, we highlight typical occurrences and burned areas rather than absolute values. As the typical number of fire occurrences may vary by department, this representation seems most effective and interpretable for firefighters. Predictions are always made daily.
Figure 3 compares class distributions between the Mediterranean basin and the rest of France. While daily fire occurrence levels are similar nationwide (except for class 4), longer horizons reveal distinct seasonality and higher fire intensity in the Mediterranean. For the burned area, class 4 shows high variance, suggesting that extreme events are scattered and department-specific. This supports the idea that our clustering method enhances predictability but may obscure moderate events when extreme cases dominate a department’s historical profile. The resulting class distributions show a strong imbalance, with class 0 dominating. Class 1 is the most frequent among positive classes for both targets, especially for the burned area. A specific histogram for each horizon is available in the supplementary materials.
Figure 3: Distribution of fire occurrence (a) and burned area (b) classes risk in the Mediterranean basin and the rest of France. Histograms have been computed relative to the category -Mediterranean (orange) or not (blue)- taking all horizons together.
3.2 Processing Features
Table 1: Summary of features used in this study. ’-’ means same as above.
Variables Frequency Source Variables Frequency Source
Meteorological Topographic
Temperature 12h, 16h Meteostat Elevation Static IGN
Dew Point - - Forest landcover - -
Precipitation - - Landcover - Cosia
Wind Direction - - NDVI, NDSI, NDMI, NDBI, NDWI 15 days GEE (landsat 1+2)
Wind Speed - - Swelling-shrinking of clays - -
Precipitation in Last 24 hours - -
Snow height - -
Sum of last 7 days rain drop - -
Day since last rain 12h -
Nesterov - firedanger
Munger - -
KBDI - -
Angstroem - -
BUI, ISI, FFMC, DMC, FWI, - -
Daily severity rating - -
Precipitation Index last 3, 5, and 9 days - Calculated
Socio-Economic Historical
Highway Static OSMNX [1] Past risk Daily Calculated
Population - Kontur Past risk burned area - -
Calendar Daily Cluster Static -
The features used in this study fall into four categories (Table 1): Meteorological, Topographic, Socio-Economic, and Historical. Most were rasterized into 2 km resolution datacubes for consistency, except for historical data, which were computed at the site level (see Figure 2).
- Meteorological data were calculated on an $11\times 1$ 11 grid per department. Fire indices followed standard methods and were reset annually. The precipitation index, based on Chen et al. [4], captures short-term rainfall variation and was computed directly on the 3D raster.
- Past risk and burned area features represent prior fire activity, processed using cubic kernel convolution (5 classes) and shifted by one day. The kernel size was based on the average fire sequence duration, with a 3-day inactivity threshold defining sequence boundaries.
- Cluster variables were derived using K-Means clustering with Dynamic Time Warping  [13], grouping departments with similar fire patterns.
- Calendar features capture daily context (e.g., weekdays, holidays, curfews) and were encoded using CatBoost [16] based on nationwide fire totals. Aggregated statistics (mean, sum, min, max) were computed across encoders to represent calendar risk.
- Categorical features were also CatBoost-encoded using fire counts per pixel. Where needed, rasters were also split by subcategories (e.g., primary vs. secondary roads) to preserve detail lost at 2 km resolution. NDVI, NDSI, NDBI, NDMI, and NDWI (Normalized Difference Vegetation, Snow, Build-up, Moisture, and Water Index) are satellite indices calculated using Landsat 1 and 2 bands. These indices represent key environmental characteristics such as vegetation density, snow cover, built-up areas, soil moisture, and surface water presence—all of which are relevant for assessing fire risk. The COSIA database originally contains 18 land cover classes, but some classes (such as swimming pools) may not be particularly useful for forest fire analysis. Therefore, we applied a logical reduction to retain only the land cover classes most relevant to forest fires.
Except for Department, Historical, and Calendar data, all spatial features were aggregated (min, max, mean) by cluster. Finally, all variables—aside from Past risk and Past Burned Area—were standardized.
3.3 Features selection
This process produced a total of 261 features. Those with no variance and those highly correlated with each other (Pearson, Spearman, and Kendall correlation threshold of 0.95) were removed, retaining those with the highest variance. This selection resulted in 179 features. The final feature list is accessible in the supplementary materials. The final dataset is composed of 2,392 unique dates (with 1,944 dates corresponding to a fire in France), across 94 departments, for a total of 224,942 rows.
4 Models
Figure 4 shows the training and evaluation process, the models used, and the architectures of the deep learning models. A detailed description of the deep learning layer parameters and tree-based configuration is available in the supplementary materials. For each model independently, we address the data imbalance problem by testing various proportions of class 0, starting from $5\%$ to $100\%$, with a step of $5\%$. We select the percentage that maximizes the Intersection over Union on the validation set. Time series models (LSTM, GRU, DilatedCNN, and GRUGNN) were tested on three time series lengths (5, 10, 15). We selected the length based on the scores on the whole dataset. The selected lengths were tested for both targets at each horizon and are available in the supplementary materials. All models were trained with CrossEntropy loss. The learning rate for deep learning models was set to 0.00005. All model parameters were tuned based on validation performance, mainly on horizon 0. The FWI system is based on the classification proposed by Dupire et. al [7]. The objective of the GRUGNN is to test whether varying temporal features and static features should be included in the same process or not. We define spatial features as those that are static (e.g., topography), while temporal features vary daily (e.g., meteorology)
Figure 4: Training and evaluation process and models architecture used in this article. Evaluation part show a comparison between FWI, LSTM, and Catboost at day 0. A full resolution image is available in supplementary materials.
We evaluate the models’ performance on each target using two main metrics:
  1. 1.
    Binary F1 score, which measures the performance of predicting the presence of at least one fire. Additionally, we report the precision and recall scores.
  2. 2.
    Intersection over Union (IoU), which measures how well the predicted risk aligns with actual risk when an event occurs. IoU is well-suited for multiclass wildfire prediction as it accounts for class uncertainty and preserves class ordinality—predicting class 1 instead of 4 is penalized less than predicting 0.
5 Results
Tables 2 and 3 present the model performance for predicting fire occurrence (FO) and burned area (BA) across four forecast lead times: $1,7,15$, and $31$ days. For each target, we report two complementary metrics: a global metric (x) computed over all departments combined, and an area-normalized score (y) computed as the area under the performance curve per department. The global metrics reflect aggregate accuracy on the entire test set, whereas the area-normalized metrics treat each department equally. This separation highlights spatial generalization by evaluating how well each model predicts fires in every department, in addition to its overall performance.
- Model Performance on global metrics
The CatBoost ensemble emerges as the top performer in the fire risk prediction benchmark, achieving the highest scores on both targets. For example, at the longest forecast horizon (31 days), CatBoost reaches an F1 score of 0.71 for fire occurrence—the best among all models—along with an intersection-over-union (IoU) of 0.45 for occurrence and 0.48 for burned area. Most other methods attain slightly lower but comparable global scores: XGBoost and even logistic regression perform nearly as well (with 31-day F1 scores around the high 0.60s), whereas deep learning models generally lag behind these simpler approaches.Longer prediction lead times consistently yield better model performance.
- Model Performance on Area-Normalized Across Departments
FWI yields higher area-normalized F1 scores across departments than the learned models. For instance, at a 31-day horizon, FWI attains an area-normalized F1 of 0.55, exceeding CatBoost’s 0.53. However, its IoU is comparatively low. This discrepancy likely stems from FWI’s lack of calibration to actual fire counts or regional conditions, which causes it to broadly overpredict risk. While such coverage yields high recall, FWI’s low overall accuracy (e.g., 0.15 global F1 at 1 day) limits its practical utility. In contrast, deep learning models often outperform tree-based models on area-normalized F1 and IoU, indicating better geographic generalization. Although area scores often correlate with recall, we observe marginal cases (e.g., GRU vs. DilatedCNN at 7 days, where both achieved comparable area-normalized scores—0.38 and 0.37—but different recall: 0.75 and 0.71, respectively), suggesting that recall alone does not fully explain generalization performance.
- Binary vs. Multi-classification
Binarizing the fire prediction task reveals clear trade-offs, with binary models achieving higher area-normalized metrics but lower global metrics such as Catboost at 15 days with 0.64, 0.5 in binary versus 0.68, 0.43 for the multi-class version—reflecting improved regional coverage. An exception exists at 15 days, with the MLP binary obtaining lower scores in both metrics. Figure 5 shows that for GRU, the multi-class model excels in high-fire departments but underperforms beyond 25 fires, where binary performance remains more stable. Similar curves for CatBoost and MLP are available in the supplementary materials. Figure 6 illustrates Catboost prediction in Bouches-du-Rhône and Charente. Multi-class predictions in Bouches-du-Rhône capture fire severity, while in Charente, binary models yield better detection despite missing fire count granularity. These results confirm that multi-class models perform best in high-activity zones, whereas binary models offer more consistent coverage in lower-risk areas. It seems that this is due to classes 2, 3, and 4 are difficult to predict due to their rarity, where classification, by assigning everything to class 1, somewhat mitigates this issue.
- Fire Occurrence vs. Burned Area
Predictions for BA consistently show better area-normalized performance than those for FO. For instance, at a 7-day forecast horizon, models achieve an IoU of approximately 0.42 for BA, compared to about 0.39 for FO, indicating higher spatial accuracy in predicting burned areas. This difference in predictability is largely attributed to class distribution: the positive class 1 (minor fires) occurs much more frequently in BA than in FO, reducing both class imbalance and prediction variance. Furthermore, the clustering approach aggregates extreme fire sizes into a single highest category. As a result, most BA predictions fall within class 0 (no fire) and class 1, leading to a narrower and more stable range of outputs. This reduced output space simplifies the prediction task, making BA significantly easier to predict than FO.
- Combining all samples’ performance and departments’ performance
The final model performance was evaluated by multiplying global accuracy (x) with area-normalized performance (y), resulting in a composite score that captures both overall effectiveness and spatial consistency. The final score is illustrated in Figure 7. For IoU on FO, GRU achieves the highest performance at 0 days, CatBoost leads at 7 and 31 days, and MLP performs best at 15 days with CatBoost being second. Models are followed by GraphCast (0 days), DilatedCNN (7 days), and LSTM (31 days). Regarding the F1 score for FO, GraphCast performs best at 0 days, while CatBoost Binary leads at 7 and 15 days. However, as the prediction horizon increases, the probability of at least one fire occurring becomes less interpretable. The highest score at 31 days is achieved by the multi-class version of CatBoost. For IoU on Burned Area (BA), GRUGNN performs best at 0 days (with Logistic Regression as runner-up), while CatBoost leads at all longer horizons, with GNN-based models (GRUGNN, GraphCast) and Logistic Regression consistently ranking among the top performers. These trends indicate that deep learning models are more effective for short-term forecasts, whereas CatBoost provides stable and high accuracy across all time horizons. When comparing GRUGNN to GRU, we found that the GNN-based model tends to perform better at longer time horizons (15–31 days) for FO, and at 31 days for BA. For short-term predictions, the GRU model remains superior. As discussed further in Section 6, this may suggest that decoupling spatial from temporal features helps the model better capture spatial patterns when they are most significant. That said, the performance differences remain relatively small, and this observation should therefore be considered preliminary.
Table 2: Model comparison for Fire Occurrence (FO) and Burned Area (BA) at 0-day and 7-days prediction horizons. “LR” stands for Logistic Regression. We do not show FWI performance on BA as it is made for fire occurrence. In addition, IoU is not shown for binary models as it is not comparable with multi-classification. The first value is the global score, and the second value (when noted) is the area-normalized score.
Model FO (0 Day) BA IoU
F1 Rec Prec IoU
FWI 0.15, 0.14 0.84 0.09 0.05, 0.05 -
LR 0.42, 0.14 0.50 0.36 0.24, 0.08 0.23, 0.08
XGBoost 0.42, 0.14 0.52 0.36 0.24, 0.08 0.24, 0.07
CatBoost 0.43, 0.13 0.52 0.37 0.24, 0.07 0.23, 0.06
MLP 0.42, 0.14 0.60 0.32 0.24, 0.08 0.23, 0.06
GRUGNN 0.41, 0.14 0.66 0.24 0.24, 0.08 0.24, 0.08
GraphCast 0.41, 0.16 0.61 0.31 0.23, 0.09 0.22, 0.07
LSTM 0.41, 0.15 0.64 0.30 0.23, 0.09 0.23, 0.07
GRU 0.43, 0.15 0.57 0.35 0.24, 0.09 0.23, 0.07
DilatedCNN 0.40, 0.16 0.54 0.32 0.22, 0.09 0.22, 0.07
CatBoost Binary 0.42, 0.15 0.63 0.31 - -
MLP Binary 0.40, 0.15 0.58 0.32 - -
GRU Binary 0.37, 0.17 0.76 0.24 - -
Model FO (7 Days) BA IoU
F1 Rec Prec IoU
FWI 0.38, 0.38 0.84 0.29 0.17, 0.17 -
LR 0.63, 0.33 0.68 0.59 0.39, 0.20 0.41, 0.20
XGBoost 0.63, 0.34 0.68 0.59 0.39, 0.20 0.42, 0.20
CatBoost 0.64, 0.35 0.69 0.59 0.40, 0.21 0.42, 0.20
MLP 0.61, 0.37 0.73 0.52 0.34, 0.21 0.41, 0.19
GRUGNN 0.59, 0.37 0.80 0.47 0.34, 0.21 0.40, 0.21
GraphCast 0.61, 0.34 0.66 0.56 0.37, 0.20 0.40, 0.20
LSTM 0.58, 0.36 0.72 0.47 0.37, 0.22 0.40, 0.19
GRU 0.60, 0.37 0.75 0.50 0.37, 0.22 0.41, 0.21
DilatedCNN 0.60, 0.37 0.71 0.52 0.38, 0.22 0.39, 0.20
CatBoost Binary 0.60, 0.40 0.69 0.59 - -
MLP Binary 0.54, 0.39 0.86 0.40 - -
GRU Binary 0.58, 0.38 0.75 0.48 - -
Table 3: Model comparison for Fire Occurrence (FO) and Burned Area (BA) at 15-days and 31-days prediction horizons.
Model FO (15 Days) BA IoU
F1 Rec Prec IoU
FWI 0.47, 0.47 0.81 0.39 0.22, 0.22 -
LR 0.67, 0.44 0.72 0.62 0.41, 0.27 0.44, 0.27
XGBoost 0.67, 0.42 0.71 0.64 0.41, 0.25 0.44, 0.26
CatBoost 0.68, 0.43 0.73 0.67 0.43, 0.26 0.45, 0.28
MLP 0.63, 0.47 0.82 0.51 0.40, 0.29 0.45, 0.25
GRUGNN 0.64, 0.45 0.80 0.53 0.40, 0.27 0.42, 0.28
GraphCast 0.63, 0.46 0.80 0.52 0.39, 0.28 0.43, 0.29
LSTM 0.63, 0.44 0.76 0.54 0.40, 0.27 0.43, 0.26
GRU 0.64, 0.45 0.76 0.56 0.40, 0.27 0.44, 0.27
DilatedCNN 0.64, 0.44 0.72 0.57 0.40, 0.26 0.42, 0.26
CatBoost Binary 0.64, 0.50 0.86 0.51 - -
MLP Binary 0.63, 0.46 0.78 0.53 - -
GRU Binary 0.61, 0.46 0.79 0.50 - -
Model FO (31 Days) BA IoU
F1 Rec Prec IoU
FWI 0.47, 0.55 0.81 0.47 0.25, 0.25 -
LR 0.69, 0.54 0.77 0.64 0.43, 0.33 0.46, 0.35
XGBoost 0.71, 0.50 0.71 0.71 0.44, 0.31 0.46, 0.33
CatBoost 0.71, 0.53 0.76 0.67 0.45, 0.34 0.48, 0.35
MLP 0.69, 0.53 0.77 0.63 0.42, 0.33 0.46, 0.34
GRUGNN 0.68, 0.53 0.81 0.59 0.43, 0.33 0.45, 0.35
GraphCast 0.67, 0.47 0.68 0.67 0.41, 0.29 0.45, 0.35
LSTM 0.68, 0.54 0.77 0.61 0.43, 0.34 0.44, 0.35
GRU 0.67, 0.52 0.76 0.61 0.43, 0.33 0.44, 0.34
DilatedCNN 0.66, 0.53 0.72 0.61 0.41, 0.33 0.41, 0.34
CatBoost Binary 0.65, 0.58 0.91 0.51 - -
MLP Binary 0.66, 0.57 0.84 0.55 - -
GRU Binary 0.67, 0.57 0.82 0.57 - -
Figure 5: Area‑normalized F1 comparison for the GRU model: multi‑class versus binary formulations as the number of departments varies.
Figure 6: Catboost prediction summary at 0 days horizons for Bouches du Rhone in multi-classification (a) and binary (b), and Charente in multi-classification (c) and binary (d).
Figure 7: Final score achieved by each model.
6 Features importance
We investigated feature importance by computing SHAP values on the multi-class CatBoost model. Figure 8 shows the top 15 features for fire occurrence across different forecast horizons (0, 7, 15, and 31 days). A corresponding figure for burned area prediction is provided in the supplementary materials.
At 0 day, fire prediction is primarily driven by history and short-term signals. Encoded temporal variables, spatial clustering, and immediate weather conditions dominate the model’s decision-making. Fire danger indices, relative humidity, and vegetation-specific factors—such as the presence of pine trees—contribute meaningfully, highlighting the importance of real-time environmental conditions. By 7 days, we found similar results, but Historical features become more influential, along with temporal indicators such as the month. By 15 days, the shift toward structural and seasonal features continues. Fire history, time-of-year encodings, and forest topographic indicators become central to the model’s predictions. By 31 days, compared to shorter-term forecasts, the model relies almost exclusively on historical, static, or slowly evolving features. Vegetation and NDVI gain importance, and socio-cultural variables also start to contribute.
Figure 8: Top 15 features computed on multi-classification Catboost model on different time horizons for fire occurrence. ID encoder corresponds to Department ID
7 Future work
In this section, we outline several directions for future work aimed at improving current model performance.
- Clustering
Current risk classes are based solely on historical fire occurrences at the departmental level. Incorporating seasonal data and clustering similar regions may help capture risk dynamics that go beyond purely historical patterns.
- Ordinal Classification Training
Risk levels are inherently ordered, yet current models treat them as independent categories. Future work could leverage ordinal-aware loss functions (e.g., Kappa loss) to exploit this structure and better handle class imbalance.
- Federated Learning
A single global model may overlook important regional patterns. Federated learning would allow each region or cluster to train a specialized model while still benefiting from shared knowledge to enhance generalization. This approach would also facilitate the inclusion of overseas departments and Corsica, which have been excluded thus far due to their significantly different vegetation and meteorological conditions compared to mainland France.
- Filtering Features
Our analysis of feature importance revealed that historical variables are strong predictors in regions with frequent past fire activity. While these features help the model capture seasonal fire trends, their static nature may bias predictions toward past patterns and overlook emerging fire-prone areas. Reducing their influence during training could help the model prioritize more transferable features. We also acknowledge that 2D convolutional networks such as ResNet and ConvLSTM were not included in our benchmark. Although effective for spatial tasks like fine-scale map prediction, their high computational cost made them unsuitable for our department-scale feature set. While they seem irrelevant at short-term forecast, spatial features for long-term forecasts gain in influence as highlighted in Section 6. We plan to explore 2D CNNs (after dimensionality reduction) in future work.
8 Conclusion
We present the first nationwide benchmark for wildfire risk prediction across all metropolitan departments in France. By analyzing this approach, we report multiple findings. A five-level ordinal classification task provides richer and more actionable information than a binary approach, particularly for high risk departments. However, multi-class models are sensitive to the clustering scheme: K-means clustering for burned area tends to compress moderate risk levels while exaggerating extreme events, making this approach currently less suitable than regression. Global versus local evaluation reveals that models may appear accurate in aggregate while underperforming in sparsely burned departments, highlighting the importance of combining overall and area-normalized metrics. CatBoost consistently performs best overall, while deep learning models tend to outperform in short-term (day-ahead) forecasts. Decoupling temporal and spatial features appears promising for enhancing spatial interpretation but requires further investigation. Feature importance analysis reveals an over-reliance on historical fire signals, which limits the model’s ability to generalize to emerging risk areas. Future work will explore tail-robust discretization strategies, ordinal-aware loss functions, federated learning, and 2D CNN architectures following dimensionality reduction, to capture effectively fine-scale spatial patterns.